OpenStudy (anonymous):

@Ganeshire8 Experimental or Theoretical Probability? When you flip a coin, there are only 2 sides - heads and tails. P(H) = 1/2 = 0.5 P(T) = 1/2 = 0.5 Experimental Theoretical Cannot be determined Neither

ganeshie8 (ganeshie8):

you're not doing any experiment - so its Theoretical

OpenStudy (anonymous):

Thanks, can you help me with one more? :)

ganeshie8 (ganeshie8):

sure, ask

OpenStudy (anonymous):

Experimental or Theoretical Probability? You toss a coin 50 times and you end up with 20 heads and 30 tails P(H) = 20/50 = 0.4 P(T) = 30/50 = 0.6 Experimental Theoretical Cannot be determined Neither

ganeshie8 (ganeshie8):

what do u think, you have conducted an experiment here - u tossed the coin several times and based on ur results you're calculating the probability... so... ?

OpenStudy (anonymous):

So would it be Experimental?

ganeshie8 (ganeshie8):

Yes !

OpenStudy (anonymous):

Yay!! Thanks so much!

ganeshie8 (ganeshie8):

yw